OpenStudy (anonymous):

Determine whether a triangle can be formed with the given side lengths. If so, use Heron’s formula to find the area of the triangle. a = 240 b = 122 c = 177 10,414.95 5897.53 5883.43 No triangle is formed.

OpenStudy (anonymous):

you get a triangle for sure

OpenStudy (anonymous):

ok could you help me find the area?

OpenStudy (anonymous):

\[A=\sqrt{s(s-a)(s-b)(s-c)}\]

OpenStudy (anonymous):

\[s=\frac{a+b+c}{2}\]

OpenStudy (anonymous):

\[\frac{240+122+177}{2}=\frac{539}{2}=269.5\]

OpenStudy (anonymous):

\[A=\sqrt{269.5(269.5-240)(269.5-122)(269.5-177)}\]

OpenStudy (anonymous):

then a calculator

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i= \sqrt{269.5%28269.5-240%29%28269.5-122%29%28269.5-177%29}

OpenStudy (anonymous):

10,414.95??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw